Operation Summary for October 19, 2011 - October 21, 2011 |
Wednesday October 19 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), with the Transfer Hall, MiniBooNE, and the Meson M03 area prepared for accesses later in the morning, and with the NuMI experiment taking beam. During the day shift, at 9:05 AM, Booster RF stations 3, 12, 14, and 19 tripped off over a 25-minute period; everything reset. An RF person reported that he was working on BRF12. Operators prepared Meson enclosures M01 and M02 for a one-hour access by Safety personnel at 9:34 AM. Operators prepared Enclosure B for a one-hour access at 9:36 AM. At 10:13 AM, operators prepared the F3 Manhole for a one-hour access by FESS personnel to repair a sump pump. At 11:05 AM, a Linac person reported problems with the low energy RF stations that was inhibiting beam; they began work on LRF3. Linac came online and Operations resumed sending beam to MiniBooNE at 10:50 AM. At 11:12 AM, Chiller #3 at CUB tripped off; a CUB person reported that they had lost a pump. The Duty Electrician investigated. Linac personnel turned off equipment when the 55 LCW system temperatures jumped up by six degrees; this LCW system gets its cooling from Chiller #3. Operations received the Operational Readiness Clearances to send beam to the Meson FTBF experiments T-1008 and T-1015. Operators brought the Antiproton Source (Pbar) online at 12:15 PM; Pbar personnel will conduct studies as soon as Linac resumes sending beam. Operators reported at 12:55 PM that the Duty Electrician had replaced the chiller #3 pump starter module at CUB. Linac personnel reported problems with LRF5 at 1:36 PM; these were resolved by 2:45 PM and Linac resumed sending beam. Due to bad beam positions, MI personnel began tuning their beam at 3:03 PM. Pbar personnel canceled their study at 3:11 PM. BRF3 and 15 tripped off at 3:50 PM; they reset. During the evening shift, Booster RF personnel reported at 4:01 PM that they were changing the BRF12 modulator. Operations established beam to NuMI and MiniBooNE at 4:09 PM. Operations established beam to T-1008 at 4:20 PM. After much tuning, MI personnel reported at 4:35 PM that they might have an aperture restriction in the 602 area; their investigations continued. Operations disabled beam to T-1008 at 11:59 PM. Thursday October 20 The midnight shift began with Operations monitoring beam to the NuMI and MiniBooNE experiments. During the midnight shift, at 12:19 AM, operators reported that the Linac Buncher gradient had started dropping; operators investigated and then called in a Linac person. At 4:18 AM, a Linac person reported that she had stabilized the gradient; this will be further investigated later in the morning. Operations halted beam and turned off MI at 7 AM in preparation for an access to investigate a possible beam aperture restriction. During the day shift, Neutron Therapy Facility (NTF) personnel reported at 8:19 AM that they would be treating one patient at 8:30 AM. A power Supply person began work on a MI Lambertson (LAM42) power supply at 9:09 AM. MI personnel reported at 9:54 AM that a bad bellows was the probable cause of the aperture restriction. Operators prepared the Transfer Hall for access at 11:37 AM. Alignment personnel in MI reported at 12:51 PM that the beam valve in front of the bellows had dropped .2 inches; when they moved the valve back to its normal position. Mechanical technicians will let up the vacuum and prepare to replace the bellows. Operators prepared the Booster for access at 1:42 PM. Operators searched and secured the Transfer Hall at 3:28 PM. Personnel working on the MI bellows found that the bellows wasn’t the problem; they stuck a camera (Bore Scope) in the beam pipe, looked downstream, and found that a ferrite rod in a microwave absorber had dropped, blocking the beam. During the evening shift, Mechanical technicians removed the ferrite rod from the microwave absorber, closed up the beam pipe, and began the vacuum pumpdown. Booster came out of access at 4:02 PM. Power Supply personnel continued to work on the Booster West Anode Power Supply; when they had finished, Operations established beam to the Booster dump at 6:15 PM. Operators reported at 8:20 PM that the MI vacuum was good enough to get the beam permits. Operators successfully conducted a MI hipot at 8:43 PM. Operations sent beam to MiniBooNE at 9:07 PM. Operations ramped MI and then began sending only a low intensity beam to NuMI at 9:12 PM; operators will increase beam intensity as vacuum improved. Sometime around 9:30 PM, Operations began sending beam to T-1008. Friday October 21 The midnight shift began with Operations sending beam to the NuMI, MiniBooNE, and T-1008 experiments. During the midnight shift, Operations disabled beam to T-1008 at 2:09 AM. Pbar personnel began a study at 6 AM. Operations increased the beam intensity to NuMI at 7:38 AM as the MI vacuum continued to improve. At 7:45 AM, NTF personnel began preparing to treat a patient later in the morning. Operators reported minor downtimes due to Linac and Booster RF trips. Operations Coordinator (OPSCO) Plans for Today and the Weekend The plans for today and the weekend are to raise the proton intensity and to allow Booster and Pbar personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 10/21/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |